Solve for "c" using the substitution method.
5x-7y=c
-15x+21y=9

Answer by ankor@dixie-net.com(22740)

Solve for "c" using the substitution method.
5x - 7y = c
-15x+21y = 9
Use the 1st equation for substitution
5x = 7y + c
Divide by 5
x = 1.4y + .2c
:
Replace x with (1.4y+.2c) in the 2nd equation
-15(1.4y + .2c) + 21y = 9
-21y - 3c + 21y = 9
-3c = 9
c = 9/-3
c = -3